40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 232     Entry time: Mon Jan 27 14:31:28 2020
Author: Maria G 
Type: Progress 
Category: BS BRDs 
Subject: BRD 7 and BRD 8 

From January 14 to the 21, Bounce BRD 7's frequency drifted -2.68% which differes from the drift logged on January 14 which was 0.15% on the opposite direction. The Roll of BRD 7's frequency drifted -4.41% which is a big change from the previous logged drift of 0.06%.

The Roll side for BRD 7 on January 21 was much closer to the target frequency. It was logged as 24.36 Hz and the target is 24.34 Hz. The Bounce side was also logged as much closer to the target frequncy. The target is 16.69 Hz and the frequency logged was 16.80 Hz.

The Q factors for both the Roll side and the Bounce side have decreased from the previous entry. Roll decreased from 108 Hz to 100 Hz and Bounce decreased from 87 Hz to 79 Hz.

BRD 8's frequency was tested for the first time. Roll side's frequency was 24.57 Hz. Bounce side's frequency was 17.956 Hz. Roll side is close to the target frequency, but Bounce side is off by 1.26 Hz.

 

The length for the new copper masses will be computed.

Attachment 1: BRD_7_v5_and_BRD_8_v5.xlsx  21 kB  Uploaded Mon Jan 27 16:31:50 2020
Attachment 2: Copper_Mass_Lengths.xlsx  10 kB  Uploaded Tue Jan 28 14:35:32 2020
ELOG V3.1.3-